# Variational Sparse Paired Autoencoders (vsPAIR) for Inverse Problems and Uncertainty Quantification

### v1

@@ B.1 MNIST Example
B.1.1 MNIST Model Details
@@ 5 Conv2D ($3\times 3$) $16\to 1$ $1\times 28\times 28$
  • Sparse variational encoder. Three linear heads produce $\mu,\log\sigma^{2},\log\omega\in\mathbb{R}^{\ell}$. Gate probabilities $\omega_{i}\in(0,1]$ are obtained via a clamped exponential $\omega=\exp(-\mathrm{ReLU}(-\log\omega))$. Samples are drawn using the hard-concrete relaxation with temperature $c=50$ and straight-through estimation (see Section 2.2).
@@
B.1.2 MNIST Hyperparameters and Training Details
@@
For the sparse models (sVAE and vsPAIR), the beta hyperprior is set to $\mathrm{Beta}(1,127)$, which has mean $\frac{1}{128}\approx 0.008$ and concentrates near zero, encouraging high sparsity (i.e., a high prior probability $\rho$ of each dimension being inactive). The hard-concrete gate temperature is $c=50$ for all sparse models. PAIR uses only MSE reconstruction losses (no KL terms), with the QoI and observation reconstruction weighted by $\lambda_{1}$, respectively $\lambda_{2}$.
B.2 CT Example
@@ 8 Linear $15,\!488\rightarrow 8,\!192$ 8,1928,\!192
Further details are outlined in Table 11.
@@ 6 Conv2D ($3\times 3$) $16\rightarrow 1$ $1\times 88\times 88$
The LPD models used adapt the implementation and networks in [19] for our down-sampled setting. All Convolutions use kernel size $3\times 3$ with stride 1, and padding 1.
B.2.2 CT Hyperparameters and Training Details
Two different sets of parameters were used to train the models showcased Figure 9 and those in Figure 10 and Figure 11. The primary difference is that sparsity was enforced strongly in the latter, while the former targeted high-fidelity reconstructions. Correspondingly, the respective pairs $(\alpha_{0},\beta_{0})$ differ between the two. All hyper- and training parameters are listed in Table 12.
@@ B.2.2 CT Hyperparameters and Training Details
Appendix C Theoretical Results
C.1 Proof of Theorem 1
The expression for mean follows from the conditional independence of $Z_{x}\mid X$ and $Y$ and the towering property of conditional expectation:
$\begin{aligned} &\mathbb{E}[Z_{x}\mid Y]=\mathbb{E}[\mathbb{E}[Z_{x}\mid Y,X]\mid Y]\\ &=\mathbb{E}[\mathbb{E}[Z_{x}\mid X]\mid Y]\quad\\ &=\mathbb{E}[BX+c\mid Y]\\ &=B\mathbb{E}[X\mid Y]+c\end{aligned}$
The expression for covariance follows similarly, but is more involved. The conditional covariance is given by
${\rm Cov}(Z_{x}\mid Y)=\mathbb{E}[(Z_{x}-\mathbb{E}[Z_{x}\mid Y])(Z_{x}-\mathbb{E}[Z_{x}\mid Y])^{\top}\mid Y]$
to which we add and subtract $\mathbb{E}[Z_{x}\mid X]$ in the following manner:
$\begin{aligned} &\mathbb{E}[(Z_{x}-\mathbb{E}[Z_{x}\mid X]+\mathbb{E}[Z_{x}\mid X]-\mathbb{E}[Z_{x}\mid Y])(Z_{x}-\mathbb{E}[Z_{x}\mid X]+\mathbb{E}[Z_{x}\mid X]-\mathbb{E}[Z_{x}\mid Y])^{\top}\mid Y]\\ &=\mathbb{E}[(Z_{x}-\mathbb{E}[Z_{x}\mid X])(Z_{x}-\mathbb{E}[Z_{x}\mid X])^{\top}\mid Y]&&{\rm(1)}\\ &+\mathbb{E}[(\mathbb{E}[Z_{x}\mid Y]-\mathbb{E}[Z_{x}\mid X])(\mathbb{E}[Z_{x}\mid Y]-\mathbb{E}[Z_{x}\mid X])^{\top}\mid Y]&&{\rm(2)}\\ &+\mathbb{E}[(Z_{x}-\mathbb{E}[Z_{x}\mid X])(\mathbb{E}[Z_{x}\mid Y]-\mathbb{E}[Z_{x}\mid X])^{\top}\mid Y]&&{\rm(3)}\\ &+\mathbb{E}[(\mathbb{E}[Z_{x}\mid Y]-\mathbb{E}[Z_{x}\mid X])(Z_{x}-\mathbb{E}[Z_{x}\mid X])^{\top}\mid Y]&&{\rm(4)}.\end{aligned}$
We now proceed line-by-line.
  1. (1)
    Using the independence of $Z_{x}\mid X$ and $Y$, we write (1) as
    $\mathbb{E}[(Z_{x}-\mathbb{E}[Z_{x}\mid X,Y])(Z_{x}-\mathbb{E}[Z_{x}\mid X,Y])^{\top}\mid Y],$
    which by the towering property we can write as
    $\mathbb{E}\!\left[\mathbb{E}\!\left[(Z_{x}-\mathbb{E}[Z_{x}\mid X,Y])(Z_{x}-\mathbb{E}[Z_{x}\mid X,Y])^{\top}\,\middle|\,X,Y\right]\,\middle|\,Y\right].$
    This is identical to
    $\mathbb{E}[{\rm Cov}(Z_{x}\mid X,Y)\mid Y],$
    which simplifies to
    $\mathbb{E}[{\rm Cov}(Z_{x}\mid X)\mid Y]=\mathbb{E}[\Sigma_{Z_{x}}\mid Y]=\Sigma_{Z_{x}}.$
  2. (2)
    We first note that
    $\mathbb{E}[Z_{x}\mid Y]=\mathbb{E}[\mathbb{E}[Z_{x}\mid X,Y]\mid Y]$
    which allows us to rewrite (2) as
    $\begin{aligned} &\mathbb{E}[(\mathbb{E}[Z_{x}\mid X]-\mathbb{E}[\mathbb{E}[Z_{x}\mid X,Y]\mid Y])(\mathbb{E}[Z_{x}\mid X]-\mathbb{E}[\mathbb{E}[Z_{x}\mid X,Y]\mid Y])^{\top}\mid Y]\\ &=\mathbb{E}[(\mathbb{E}[Z_{x}\mid X]-\mathbb{E}[\mathbb{E}[Z_{x}\mid X]\mid Y])(\mathbb{E}[Z_{x}\mid X]-\mathbb{E}[\mathbb{E}[Z_{x}\mid X]\mid Y])^{\top}\mid Y]\end{aligned}$
    using the conditional independence assumption. Note that this is exactly ${\rm Cov}(\mathbb{E}[Z_{x}\mid X]\mid Y)$. In particular, $\mathbb{E}[Z_{x}\mid X]=BX+c$ which is an affine transform of $X$. Hence,
    $\begin{aligned} &{\rm Cov}(\mathbb{E}[Z_{x}\mid X]\mid Y)\\ &={\rm Cov}(BX+c\mid Y)\\ &=B{\rm Cov}(X\mid Y)B^{\top}.\end{aligned}$
  3. (3)
    We rewrite again applying the tower property:
    $\begin{aligned} &\mathbb{E}[(Z_{x}-\mathbb{E}[Z_{x}\mid X])(\mathbb{E}[Z_{x}\mid X]-\mathbb{E}[Z_{x}\mid Y])^{\top}\mid Y]\\ &=\mathbb{E}[\mathbb{E}[(Z_{x}-\mathbb{E}[Z_{x}\mid X])(\mathbb{E}[Z_{x}\mid X]-\mathbb{E}[Z_{x}\mid Y])^{\top}\mid X,Y]\mid Y].\end{aligned}$
    Note that in the inner-expectation, $\mathbb{E}[Z_{x}\mid X]-\mathbb{E}[Z_{x}\mid Y]$ is a deterministic function of $X,Y$, thus the conditioning on $X$ and $Y$ reduces our expression to
    $\begin{aligned} \mathbb{E}[\mathbb{E}[(Z_{x}-\mathbb{E}[Z_{x}\mid X])\mid X,Y](\mathbb{E}[Z_{x}\mid X]-\mathbb{E}[Z_{x}\mid Y])^{\top}\mid Y].\end{aligned}$
    In particular,
    $\mathbb{E}[(Z_{x}-\mathbb{E}[Z_{x}\mid X])\mid X,Y]=\mathbb{E}[(Z_{x}-\mathbb{E}[Z_{x}\mid X])\mid X]=0$
    using conditional independence, and the definition of conditional expectation.
  4. (4)
    This expression is symmetric to (3) and also simplifies to 0.
Together,
$\mathbb{E}_{Z_{x}\mid Y}[Z_{x}\mid Y]=B\mathbb{E}[X\mid Y]+c$
and
${\rm Cov}(Z_{x}\mid Y)=\Sigma_{Z_{x}}+B{\rm Cov}(X\mid Y)B^{\top}.$
This is clearly symmetric positive definite and assumption 2 guarantees continuity.
C.2 Proof of Theorem 2
Let $Y=AX+\varepsilon$, with $\varepsilon\sim\mathcal{N}(\varepsilon\mid 0,\Sigma)$ and $p(x\mid z_{x})=\mathcal{N}(x\mid\mu,\sigma)$. Writing $p(y\mid z_{x})$ as the marginal over $x$, we obtain
$p(y\mid z_{x})=\int_{X}p(y,x\mid z_{x})dx$ (27)
$p(y\mid z_{x})=\int_{X}p(y\mid x,z_{x})p(x\mid z_{x})dx$ (28)
$p(y\mid z_{x})=\int_{X}p(y\mid x)p(x\mid z_{x})dx$ (29)
by our conditional independence assumption. Note that this matches the marginal likelihood computation for Bayesian linear regression with a Bayesian prior on parameters. This integral can be computed in closed form, and results in a Gaussian distribution [8, 42]. The result in Theorem 2 is obtained by replacing $p(x\mid z_{x})$ by $q_{\theta_{x}}(x\mid z_{x})$.

### v2

@@ D.1 MNIST Example
D.1.1 MNIST Model Details
@@ 5 Conv2D ($3\times 3$) $16\to 1$ $1\times 28\times 28$
  • Sparse variational encoder. Three linear heads produce $\mu,\log\sigma^{2},\log\omega\in\mathbb{R}^{\ell}$. Probabilities $\omega_{i}\in(0,1]$ are obtained via a clamped exponential $\omega=\exp(-\mathrm{ReLU}(-\log\omega))$. Samples are drawn using the Bernoulli relaxation with temperature $c=50$ and straight-through estimation (see Appendix A).
@@
D.1.2 MNIST Hyperparameters and Training Details
@@
For the sparse models (sVAE and vsPAIR), the beta hyperprior is set to $\mathrm{Beta}(1,127)$, which has mean $\frac{1}{128}\approx 0.008$ and concentrates near zero, encouraging high sparsity (i.e., a high prior probability $\rho$ of each dimension being inactive). The Bernoulli approximation temperature is $c=50$ for all sparse models. PAIR uses only MSE reconstruction losses (no KL terms), with the QoI and observation reconstruction weighted by $\lambda_{1}$, respectively $\lambda_{2}$.
D.2 CT Example
@@ 8 Linear $15,\!488\rightarrow 8,\!192$ 8,1928,\!192
Further details are outlined in Table 11. All Convolutions use kernel size $3\times 3$ with stride 1, and padding 1.
@@ 6 Conv2D ($3\times 3$) $16\rightarrow 1$ $1\times 88\times 88$
The LPD models used adapt the implementation and networks in [20] for our down-sampled setting; LPD models were trained for 500 epochs with learning rate 1e-4 and batchsize 2.
D.2.2 CT Hyperparameters and Training Details
Two different sets of parameters were used to train the models showcased Figure 9 and those in Figure 10 and Figure 11. The primary difference is that sparsity was enforced strongly in the latter, while the former targeted high-fidelity reconstructions. Correspondingly, the respective pairs $(\alpha_{0},\beta_{0})$ differ between the two. All hyper- and training parameters are listed in Table 12. The KL weights are notably lower than the MNIST experiment, since we focus on the fidelity of the reconstruction. We use a batch size of 2 for all training.